# Fixed Points and Cycles of Parking Functions

### v1

@@ Abstract.
A parking function of length $n$ is a sequence $\pi=(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq i$ for $1\leq i\leq n$. In this paper we obtain some exact results on the number of fixed points and cycles of parking functions. Our proofs will be based on generalizations of Pollak’s argument. Extensions of our techniques are discussed.
@@
Parking function; Pollak’s circle argument
@@
M. Yin was supported by the University of Denver’s Professional Research Opportunities for Faculty Fund 80369-145601.
@@ 1. Introduction
Parking functions were introduced by Konheim and Weiss [2] in the study of the linear probes of random hashing functions. In the classical parking function scenario, we have $n$ parking spaces on a one-way street, labelled $1,2,\dots,n$ in consecutive order as we drive down the street. There are $n$ cars $C_{1},\dots,C_{n}$. Each car $C_{i}$ has a preferred space $1\leq\pi_{i}\leq n$. The cars drive down the street one at a time in the order $C_{1},\dots,C_{n}$. The car $C_{i}$ drives immediately to space $i$ and then parks in the first available space. Thus if $i$ is empty, then $C_{i}$ parks there; otherwise $C_{i}$ next goes to space $i+1$, etc. If all cars are able to park, then the sequence $\pi=(\pi_{1},\dots,\pi_{n})$ is called a parking function of length $n$. It is well-known and easy to see that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\pi$ is a parking function if and only if $\lambda_{i}\leq i$ for $1\leq i\leq n$. In particular, any permutation of a parking function is a parking function. Write $\operatorname{PF}_{n}$ for the set of parking functions of length $n$.
@@
The first significant result on parking functions, due to Pyke [4] in another context and then to Konheim and Weiss [2], is that the number of parking functions of length $n$ is equal to $(n+1)^{n-1}$. A famous combinatorial proof was given by Pollak (unpublished but recounted in [1] and [5]). It boils down to the following easily verified statement: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n+1]^{n}$ with componentwise addition modulo $n+1$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one parking function.
We will be interested in deriving some exact formulas for the number of fixed points and cycles of parking functions in this paper. For a parking function $\pi=(\pi_{1},\dots,\pi_{n})$, we say $i$ is a fixed point of $\pi$ if $\pi_{i}=i$. More generally, for $m\geq 1$, we say $(i_{1},\dots,i_{m})$ where the $i_{j}$’s are all distinct is an $m$-cycle of $\pi$ if $\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}$, so fixed points are simply $1$-cycles. Our proofs will be based on generalizations of Pollak’s argument. In Section 2 we study classical parking functions, in Section 3 we study prime parking functions, and in Section 4 we discuss extensions of our techniques to $(r,k)$-parking functions.
@@ 2. Classical parking functions
$\frac{1}{(n+1)^{2}}\binom{n+1}{k}\left(n^{n-k+1}+(-1)^{n+k}\right).$
@@
For $1\leq i\leq n$, let $A_{i}=|\{\pi\in\operatorname{PF}_{n}:\pi_{i}=i\}|$. By De Morgan’s law and the inclusion-exclusion principle, the number of parking functions of length $n$ with $k$ fixed points is
$\hskip 56.9055pt\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\left|\left(\bigcap_{j=1}^{k}A_{i_{j}}\right)\bigcap_{i\neq i_{j}\,\forall 1\leq j\leq k}A_{i}^{c}\right|$
@@
$=$ $\sum_{\ell=0}^{n-k}(-1)^{\ell}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k+\ell}\\ \left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\}\end{subarray}\right\}\right|,$
where the last equality follows by symmetry of parking coordinates, and by
$\left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\},$
we mean equality as a set.
To obtain a simplified expression for this last equation, we use an extension of Pollak’s circle argument. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $k+\ell$ spots for the first $k+\ell$ cars. Note that it matters whether these spots correspond to $i_{j}$ for $1\leq j\leq k$ or to $s_{t}$ for $1\leq t\leq\ell$, and altogether this can be done in $\binom{n+1}{k,\ell,n-k-\ell+1}$ ways. Then for the remaining $n-k-\ell$ cars, there are $(n+1)^{n-k-\ell}$ possible preference sequences. Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
$\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k+\ell}\\ \left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\}\end{subarray}\right\}\right|$
$\hskip 56.9055pt=\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}.$
@@
$\sum_{\ell=0}^{n-k}(-1)^{\ell}\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}=\binom{n+1}{k}\frac{((n+1)-1)^{n-k+1}+(-1)^{n+k}}{(n+1)^{2}}.$
@@
We still apply De Morgan’s law, the inclusion-exclusion principle, and an extension of Pollak’s circle argument. The major difference is that instead of selecting $k+\ell$ spots on a circle with $n+1$ spots, we select $k+\ell$ $m$-tuples on the circle. We also include an additional scalar factor $(m-1)!$ for each such tuple, which accounts for the number of ways of arranging an $m$-cycle with distinct entries. ∎
@@
From Theorem 2.3, we may readily derive that for fixed $m$, as $n$ gets large, the expected number of $m$-cycles in a random parking function is asymptotically $1/m$. This asymptotic result was obtained earlier by Paguyo [3] using other means.
@@
To obtain a simplified expression for this last equation, we use an extension of Pollak’s circle argument as in the proof of Theorem 2.1. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $m$ spots for the first $m$ cars, which can be done in $\binom{n+1}{m}$ ways. Then for the remaining $n-m$ cars, there are $(n+1)^{n-m}$ possible preference sequences. Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
|$\left|\left\{\pi\in\operatorname{PF}_{n}:\pi_{1}<\pi_{2}<\cdots<\pi_{m}\right\}\right|=\binom{n+1}{m}(n+1)^{n-m-1}.$

### v3

@@ Abstract.
A parking function of length $n$ is a sequence $\pi=(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq i$ for $1\leq i\leq n$. The index $i$ is a fixed point of the parking function $\pi$ if $\pi_{i}=i$. More generally, for $m\geq 1$, the indices $(i_{1},\dots,i_{m})$ where the $i_{j}$’s are all distinct constitute an $m$-cycle of the parking function $\pi$ if $\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}$. In this paper we obtain some exact results on the number of fixed points and cycles of parking functions. Our derivations are based on generalizations of Pollak’s argument and the symmetry of parking coordinates. Extensions of our techniques are discussed.
@@
Parking functions; Pollak’s circle argument; Fixed points and cycles
@@
M. Yin was supported by the University of Denver’s Professional Research Opportunities for Faculty Fund 80369-145601 and Simons Foundation Grant MPS-TSM-00007227.
@@ 1. Introduction
Parking functions were introduced by Konheim and Weiss [9] in the study of the linear probes of random hashing functions. In the classical parking function scenario, we have $n$ parking spaces on a one-way street, labelled $1,2,\dots,n$ in consecutive order as we drive down the street. There are $n$ cars $C_{1},\dots,C_{n}$. Each car $C_{i}$ has a preferred space $1\leq\pi_{i}\leq n$. The cars drive down the street one at a time in order $C_{1},\dots,C_{n}$. The car $C_{i}$ drives immediately to space $\pi_{i}$ and then parks in the first available space. Thus if $\pi_{i}$ is empty, then $C_{i}$ parks there; otherwise $C_{i}$ next goes to space $\pi_{i}+1$, $\pi_{i}+2$, etc., until it finds an available space to park in (if no such space exists, then $C_{i}$ leaves the street unparked). If all cars are able to park, then the sequence $\pi=(\pi_{1},\dots,\pi_{n})$ is called a parking function of length $n$. It is well-known and easy to see that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\pi$ is a parking function if and only if $\lambda_{i}\leq i$ for $1\leq i\leq n$. Equivalently, $\pi$ is a parking function if and only if
#$\#\{k:\pi_{k}\leq i\}\geq i,\hskip 5.69046pt\forall i=1,\dots,n.$
This implies that parking functions are invariant under the action of the symmetric group $\operatorname{\mathfrak{S}}_{n}$ permuting the $n$ cars, that is, permuting the list of preferences $\pi$. Write $\operatorname{PF}_{n}$ for the set of parking functions of length $n$.
@@
The first significant result on parking functions, due to Pyke [11] in another context and then to Konheim and Weiss [9], is that the number of parking functions of length $n$ is equal to $(n+1)^{n-1}$. A famous combinatorial proof was given by Pollak (unpublished but recounted in [7] and [12]). It boils down to the following easily verified statement: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n+1]^{n}$ with componentwise addition modulo $n+1$ (where $[n]$ is the standard shorthand for $\{1,\dots,n\}$). Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one parking function.
Nowadays the study of parking functions has found many applications in combinatorics, probability, computer science, and beyond. Some recent work may be found in [1, 3, 6, 8, 18]. We refer to Yan [17] for a comprehensive survey and to Carlson et al. [2] for the many generalizations of the classical parking function scenario. In this paper we explore the cycle structure of parking functions. Unlike the cycle structure of permutations which has been well studied, not much is known about the cycle structure of parking functions, especially in terms of exact enumerative results. A traditional direction is to standardize a parking function to a permutation and then take the usual cycle structure. This was discussed in detail in Stanley [15] supplementary problems for Chapter 7, utilizing the powerful parking function symmetric function machinery. Our investigation makes an effort in another direction. See Diaconis and Hicks [4] for some open problems and Paguyo [10] for some partial answers in this direction. For a parking function $\pi=(\pi_{1},\dots,\pi_{n})$, we say $i$ is a fixed point of $\pi$ if $\pi_{i}=i$. More generally, for $m\geq 1$, we say $(i_{1},\dots,i_{m})$ where the $i_{j}$’s are all distinct is an $m$-cycle of $\pi$ if $\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}$. Fixed points are simply $1$-cycles and cyclic points are points that lie on some $m$-cycle of $\pi$ where $m\geq 1$.
We are interested in deriving some exact formulas for the number of fixed points and cycles of parking functions. Our counting formulas are based on generalizations of Pollak’s argument and the symmetry of parking coordinates. In Section 2 we study classical parking functions, in Section 3 we study prime parking functions, and in Section 4 we discuss extensions of our techniques to $(r,k)$-parking functions (see Section 4.1) and present connections of our current investigation to other research areas through OEIS entries (see Section 4.2). While some OEIS entries are easy to explain, many entries remain mysterious, waiting to be explored deeper in the future.
Theorems 2.1 and 2.2 come as a particular surprise:
Theorem 2.1.
$\frac{1}{(n+1)^{2}}\binom{n+1}{k}\left(n^{n-k+1}-(-1)^{n-k+1}\right).$
Theorem 2.2.
The generating function for parking functions of length $n$ with respect to the number of fixed points is given by
$\frac{1}{(n+1)^{2}}\left((q+n)^{n+1}-(q-1)^{n+1}\right).$
In contrast, there is not such a nice and clean formula for permutations, which despite their simpler structure are characterized by the intricate generating function
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}\frac{F(k,n)z^{n}q^{k}}{n!}=\frac{\exp((q-1)z)}{1-z},$
where $F(k,n)$ is the number of permutations of $[n]$ with $k$ fixed points.
Theorems 3.1 and 3.2 are the corresponding prime versions for parking functions:
Theorem 3.1.
$\binom{n-1}{k}(n-2)^{n-k-1}.$
Theorem 3.2.
The generating function for prime parking functions of length $n$ with respect to the number of fixed points is given by $(q+n-2)^{n-1}$.
Extending the notion of fixed points to cycles, we obtain more general results for cycles in classical parking functions in Theorem 2.3 and prime parking functions in Theorem 3.3. Our exact formulas for fixed points and cycles furthermore provide insight into the asymptotic features of the cycle structure of random parking functions, particularly in connection with random functions and random permutations whose cycle structures have already been quite well understood.
  • A random parking function is a random function in $\mathcal{F}_{n}=\{f:[n]\rightarrow[n]\}$ conditioned on being in $\operatorname{PF}_{n}$.
  • At the other end of the spectrum, we have random permutations, which are special cases of random parking functions where all cars have different preferred spaces.
As stated by Diaconis and Hicks [4], in analogy with equivalence of ensembles in statistical mechanics, it is natural to expect that for some features, the distribution of the features in the “micro-canonical ensemble” ($\operatorname{PF}_{n}$) should be close to the features in the “canonical ensemble” ($\mathcal{F}_{n}$). We will show in Remark 2.5 that the expected number of $m$-cycles in a random parking function is asymptotically $1/m$ for any $m$. This coincides with the corresponding asymptotic result for random functions as well as random permutations. Nevertheless error terms from $1/m$ are not the same or uniform for the three different random objects.
  • For random permutations in the symmetric group $\operatorname{\mathfrak{S}}_{n}$, the asymptotic formula $1/m$ for the expected number of $m$-cycles is also exact. Based on this fact, Shepp and Lloyd [13] employed a generating function approach to show that the total number of cycles in a uniformly random permutation from $\operatorname{\mathfrak{S}}_{n}$ is asymptotically normal with mean and variance $\log n$.
  • For random functions in $\mathcal{F}_{n}$, the exact formula for the expected number of $m$-cycles is $(n)_{m}/(mn^{m})$, where $(n)_{m}$ is a falling factorial. Building upon this, Flajolet and Odlyzko [5] applied generating function techniques to show that the total number of cycles in a uniformly random function from $\mathcal{F}_{n}$ is asymptotically normal with mean and variance $\frac{1}{2}\log n$. Note that asymptotically, random functions have about half as many total cycles as random permutations.
  • Now comes our contribution. In Theorem 2.4, we present the exact formula for the expected number of $m$-cycles in random parking functions in $\operatorname{PF}_{n}$, and it should be tractable to use this result to check against the equivalence of ensembles heuristic. Theorem 3.4 provides the corresponding prime version.
@@ 2. Classical parking functions
$\frac{1}{(n+1)^{2}}\binom{n+1}{k}\left(n^{n-k+1}-(-1)^{n-k+1}\right).$
@@
For $1\leq i\leq n$, let $A_{i}=|\{\pi\in\operatorname{PF}_{n}:\pi_{i}=i\}|$. Recall that the number of parking functions of length $n$ is $|\operatorname{PF}_{n}|=(n+1)^{n-1}$. For temporarily fixed $1\leq i_{1}<\cdots<i_{k}\leq n$, by De Morgan’s law and the inclusion-exclusion principle,
$\left|\bigcap_{i\neq i_{j}\,\forall 1\leq j\leq k}A_{i}^{c}\right|=\left|\operatorname{PF}_{n}-\bigcup_{i\neq i_{j}\,\forall 1\leq j\leq k}A_{i}\right|$
$=$ $|\operatorname{PF}_{n}|-\sum_{\ell=1}^{n-k}(-1)^{\ell+1}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\bigcap_{t=1}^{\ell}A_{s_{t}}\right|$
$=$ $\sum_{\ell=0}^{n-k}(-1)^{\ell}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\bigcap_{t=1}^{\ell}A_{s_{t}}\right|.$
The number of parking functions of length $n$ with $k$ fixed points is thus
$\hskip 56.9055pt\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\left|\left(\bigcap_{j=1}^{k}A_{i_{j}}\right)\bigcap\left(\bigcap_{i\neq i_{j}\,\forall 1\leq j\leq k}A_{i}^{c}\right)\right|$
@@
$=$ $\sum_{\ell=0}^{n-k}(-1)^{\ell}\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k},\pi_{k+1}<\cdots<\pi_{k+\ell}\\ \pi_{1},\ldots,\pi_{k+\ell}\text{ are pairwise distinct}\end{subarray}\right\}\right|,$
where the last equality follows by symmetry of parking coordinates, since any permutation of a parking function is a parking function.
For ease of notation, let
$A_{k+\ell}=\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k},\pi_{k+1}<\cdots<\pi_{k+\ell}\\ \pi_{1},\ldots,\pi_{k+\ell}\text{ are pairwise distinct}\end{subarray}\right\}.$
Further let
$A^{\prime}_{k+\ell}=\left\{\pi\in\operatorname{PF}_{n}:\pi_{1},\pi_{2},\dots,\pi_{k+\ell}\text{ are pairwise distinct}\right\},$
$B^{\prime}_{k+\ell}=\left\{f:[n]\rightarrow[n+1]:f(1),f(2),\dots,f(k+\ell)\text{ are pairwise distinct}\right\}.$
It is clear that $|A^{\prime}_{k+\ell}|=k!\ell!|A_{k+\ell}|$. To obtain a simplified expression for $A_{k+\ell}$, we use an extension of Pollak’s circle argument. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $k+\ell$ spots for the first $k+\ell$ cars, which can be done in $\binom{n+1}{k+\ell}$ ways. Then for the remaining $n-k-\ell$ cars, there are $(n+1)^{n-k-\ell}$ possible preference sequences. Hence
$|B^{\prime}_{k+\ell}|=(k+\ell)!\binom{n+1}{k+\ell}(n+1)^{n-k-\ell}.$
Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
$|A^{\prime}_{k+\ell}|=\frac{1}{n+1}|B^{\prime}_{k+\ell}|=(k+\ell)!\binom{n+1}{k+\ell}(n+1)^{n-k-\ell-1},$
and so
$|A_{k+\ell}|=\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}.$
@@
$\sum_{\ell=0}^{n-k}(-1)^{\ell}\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}$
$=$ $\binom{n+1}{k}\sum_{\ell=0}^{n-k}(-1)^{\ell}\binom{n-k+1}{\ell}(n+1)^{n-k-\ell-1}$
$=$ $\binom{n+1}{k}\frac{1}{(n+1)^{2}}\left(\sum_{\ell=0}^{n-k+1}(-1)^{\ell}\binom{n-k+1}{\ell}(n+1)^{n-k+1-\ell}-(-1)^{n-k+1}\right)$
$=$ $\binom{n+1}{k}\frac{((n+1)-1)^{n-k+1}-(-1)^{n-k+1}}{(n+1)^{2}}.$
Theorem 2.2.
The generating function for parking functions of length $n$ with respect to the number of fixed points is given by
$\frac{1}{(n+1)^{2}}\left((q+n)^{n+1}-(q-1)^{n+1}\right).$
Proof.
Utilizing Theorem 2.1, we perform generating function calculations.
$\sum_{k=0}^{n}q^{k}\frac{1}{(n+1)^{2}}\binom{n+1}{k}\left(n^{n-k+1}-(-1)^{n-k+1}\right)$
$=$ $\frac{1}{(n+1)^{2}}\left(\sum_{k=0}^{n+1}\binom{n+1}{k}q^{k}n^{n-k+1}-\sum_{k=0}^{n+1}\binom{n+1}{k}q^{k}(-1)^{n-k+1}\right)$
$=$ $\frac{1}{(n+1)^{2}}\left((q+n)^{n+1}-(q-1)^{n+1}\right).$
@@
We still apply De Morgan’s law and the inclusion-exclusion principle. In analogy with Theorem 2.1, we need to compute
$\sum_{\ell:(k+\ell)m\leq n}(-1)^{\ell}\left((m-1)!\right)^{k+\ell}\sum_{(i^{j}_{1},\dots,i^{j}_{m}):1\leq j\leq k}\sum_{(s^{t}_{1},\dots,s^{t}_{m}):1\leq t\leq\ell}\left|\left(\bigcap_{j=1}^{k}A_{(i^{j}_{1},\dots,i^{j}_{m})}\right)\bigcap\left(\bigcap_{t=1}^{\ell}A_{(s^{t}_{1},\dots,s^{t}_{m})}\right)\right|,$
where the middle sum is over all possible $k$ $m$-cycles and the last sum is over all possible $\ell$ $m$-cycles, and all the $k+\ell$ $m$-cycles are pairwise non-overlapping. To avoid overcounting, we arrange the $m$-cycles $(i^{j}_{1},\dots,i^{j}_{m})$ so that for each $1\leq j\leq k$, $i^{j}_{1}<\cdots<i^{j}_{m}$ and further $i^{1}_{1}<\cdots<i^{k}_{1}$, and similarly for each $1\leq t\leq\ell$, $s^{t}_{1}<\cdots<s^{t}_{m}$ and further $s^{1}_{1}<\cdots<s^{\ell}_{1}$. This hence introduces an additional scalar factor $(m-1)!$ for each such $m$-cycle, which accounts for the number of ways of arranging an $m$-cycle with distinct entries.
Next we use symmetry of parking coordinates and apply an extension of Pollak’s circle argument. The major difference from Theorem 2.1 is that instead of selecting $k+\ell$ spots on a circle with $n+1$ spots, we select $k+\ell$ $m$-cycles on the circle. Again in analogy with Theorem 2.1,
$\sum_{(i^{j}_{1},\dots,i^{j}_{m}):1\leq j\leq k}\sum_{(s^{t}_{1},\dots,s^{t}_{m}):1\leq t\leq\ell}\left|\left(\bigcap_{j=1}^{k}A_{(i^{j}_{1},\dots,i^{j}_{m})}\right)\bigcap\left(\bigcap_{t=1}^{\ell}A_{(s^{t}_{1},\dots,s^{t}_{m})}\right)\right|$
$=$ |$\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi^{1}_{1}<\cdots<\pi^{k}_{1},\pi^{k+1}_{1}<\cdots<\pi^{k+\ell}_{1}\\ (\pi^{1}_{1},\dots,\pi^{1}_{m}),\dots,(\pi^{(k+\ell)}_{1},\dots,\pi^{(k+\ell)}_{m})\text{ are pairwise non-overlapping}\end{subarray}\right\}\right|$
[…]
$=$ $\frac{1}{k!\ell!}\binom{n+1}{\underbracket{m,\cdots,m}_{k+\ell\hskip 2.84544pt\text{$m$'s}},n-(k+\ell)m+1}(n+1)^{n-(k+\ell)m-1}.$
Our conclusion follows. ∎
@@
From Theorem 2.4, we may readily derive that for fixed $m$, as $n$ gets large, the expected number of $m$-cycles in a random parking function is asymptotically $1/m$. This asymptotic result was obtained earlier by Paguyo [10] using other techniques.
@@
For ease of notation, let
$A_{m}=\left\{\pi\in\operatorname{PF}_{n}:\pi_{1}<\pi_{2}<\cdots<\pi_{m}\right\}.$
Further let
$A^{\prime}_{m}=\left\{\pi\in\operatorname{PF}_{n}:\pi_{1},\pi_{2},\dots,\pi_{m}\text{ are pairwise distinct}\right\}$
and
$B^{\prime}_{m}=\left\{f:[n]\rightarrow[n+1]:f(1),f(2),\dots,f(m)\text{ are pairwise distinct}\right\}.$
It is clear that $|A^{\prime}_{m}|=m!|A_{m}|$. To obtain a simplified expression for $A_{m}$, we use an extension of Pollak’s circle argument as in the proof of Theorem 2.1. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $m$ spots for the first $m$ cars, which can be done in $\binom{n+1}{m}$ ways. Then for the remaining $n-m$ cars, there are $(n+1)^{n-m}$ possible preference sequences. Hence
$|B^{\prime}_{m}|=m!\binom{n+1}{m}(n+1)^{n-m}.$
Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
$|A^{\prime}_{m}|=\frac{1}{n+1}|B^{\prime}_{m}|=m!\binom{n+1}{m}(n+1)^{n-m-1},$
and so
$|A_{m}|=\binom{n+1}{m}(n+1)^{n-m-1}.$
@@
Corollary 2.6.
Let $\pi\in\operatorname{PF}_{n}$ be a parking function chosen uniformly at random. The expected number of cyclic points is given by
$\sum_{m=1}^{n}m!\binom{n+1}{m}/(n+1)^{m}.$
Proof.
As stated in the introduction, cyclic points are points that lie on some $m$-cycle of $\pi$ where $m\geq 1$. We apply Theorem 2.4 and note that every $m$-cycle contains $m$ cyclic points. ∎